OpenStudy (anonymous):

3 --% ×75 = _____ 4 Select one of the options below as your answer:A. .255 B. .5625 C. 2.55 D. 5.625

OpenStudy (freckles):

@michaeljordan don't repost. Please work with me to get the answer. I'm not just going to give you the answer.

OpenStudy (freckles):

So again \[\%=\frac{1}{100}\] So we have \[\frac{3}{4} \cdot \frac{1}{100} \times 75\] Which can be written as \[\frac{3}{4} \cdot \frac{75}{100}\] What does 75/100 reduce to?

OpenStudy (anonymous):

.75

OpenStudy (freckles):

Ok yes you can write 75/100 as .75 and you can also right 3/4 as .75 since 75/100 reduces to 3/4 which means 75/100=3/4 so we need to find the product of .75 and .75

OpenStudy (freckles):